Question 1 : Consider the equilibrium of AgCl(s) in water. What is the effect of adding KCl?
The reaction goes to the left
Silver chloride solubility increases
The KCl has no effect
The reaction goes to the right

Question 2 : A Brønsted-Lowry acid is defined as a substance that ______ .
acts as a proton acceptor
decreases [H3O+] when dissolved in water
increases [OH– ] when dissolved in water
acts as a proton donor

Question 3 : The pH of a solution is 4.6; what is its pOH?
9.4
4.6
11.6
2.4
1
Expert's answer
2017-04-04T11:17:07-0400
Answer: answer is “The reaction goes to the left” for reaction AgCl (s) = Ag + (aq) + Cl - (aq) and answer
is “The reaction goes to the right” for reaction Ag + (aq) + Cl - (aq) = AgCl (s)


acts as a proton donor


9.4